I'm stuck and need help! I have spent three long hours trying to figure out how they reduced a certain part of Quadratic Equation application problem. This is a part I must reduce before I can put the equation into Quadratic Formula Form. Once I get past this, the rest will not be a problem. Here is the part of the problem: X[(110/Xsquared)/600], and it reduces to: 121/6X. I can't firue out how it reduces to 121/6X. This part was derived from the formula where the cost of fuel for a van is C=Vsquared/600 where V is the speed in miles per hour. The driver is paid $5.00 per hour. Find the speed if the cost for wages and fuel for a 110-mile trip is $20.39. Can you help me?
Thanks,
Ed Smith

Answer by stanbon(75887)

This part was derived from the formula where the cost of fuel for a van is C=Vsquared/600 where V is the speed in miles per hour. The driver is paid $5.00 per hour. Find the speed if the cost for wages and fuel for a 110-mile trip is $20.39. Can you help me?
---------
C=V^2/600
20.39=V^2/600
V^2 = 600*20.39= 12234 mph
V=110.607 mph
-------
That's awfully fast for a van. Are you sure you posted everything
properly. Why did you list $5 per hour? ; it has nothing to do with the
problem as you have stated it.
-------------
Post it again if there is more information or different information.
====================
----------
Cheers,
Stan H.